The risk-free rate is 3.96% and the market risk premium is 9.00%. A stock with a β of 1.19 just paid a dividend of $2.84. The dividend is expected to grow at 20.37% for three years and then grow at 4.84% forever. What is the value of the stock?

Submit

Answer format: Currency: Round to: 2 decimal places.

unanswered

not_submitted

#3

The risk-free rate is 3.70% and the market risk premium is 7.42%. A stock with a β of 1.32 just paid a dividend of $1.09. The dividend is expected to grow at 21.86% for five years and then grow at 4.23% forever. What is the value of the stock?

Submit

Answer format: Currency: Round to: 2 decimal places.

unanswered

not_submitted

#4

Caspian Sea Drinks needs to raise $24.00 million by issuing additional shares of stock. If the market estimates CSD will pay a dividend of $2.42 next year, which will grow at 4.46% forever and the cost of equity to be 10.81%, then how many shares of stock must CSD sell?

Submit

Answer format: Number: Round to: 0 decimal places.

unanswered

not_submitted

#5

Suppose the risk-free rate is 2.27% and an analyst assumes a market risk premium of 7.89%. Firm A just paid a dividend of $1.14 per share. The analyst estimates the β of Firm A to be 1.45 and estimates the dividend growth rate to be 4.44% forever. Firm A has 250.00 million shares outstanding. Firm B just paid a dividend of $1.77 per share. The analyst estimates the β of Firm B to be 0.82 and believes that dividends will grow at 2.04% forever. Firm B has 180.00 million shares outstanding. What is the value of Firm A?

Submit

Answer format: Currency: Round to: 2 decimal places.

unanswered

not_submitted

#6

Suppose the risk-free rate is 1.25% and an analyst assumes a market risk premium of 6.18%. Firm A just paid a dividend of $1.18 per share. The analyst estimates the β of Firm A to be 1.50 and estimates the dividend growth rate to be 4.99% forever. Firm A has 284.00 million shares outstanding. Firm B just paid a dividend of $1.72 per share. The analyst estimates the β of Firm B to be 0.89 and believes that dividends will grow at 2.85% forever. Firm B has 192.00 million shares outstanding. What is the value of Firm B?
Answer #1

1.
required return=3.96%+1.19*9%=14.67000%

Price=2.84*(1.2037/1.1467)+2.84*(1.2037/1.1467)^2+2.84*(1.2037/1.1467)^3+2.84*(1.2037/1.1467)^3*1.0484/(14.67%-4.84%)
=44.43

2.
required return=3.70%+1.32*7.42%=13.49440%

Price=1.09*(1.2186/1.134944)+1.09*(1.2186/1.134944)^2+1.09*(1.2186/1.134944)^3++1.09*(1.2186/1.134944)^4+1.09*(1.2186/1.134944)^5+1.09*(1.2186/1.134944)^5*1.0423/(13.4944%-4.23%)
=24.28

3.
Number of shares=24000000/(2.42/(10.81%-4.46%))=629752.07

4.
=250*1.14*1.044/(2.27%+1.45*7.89%-4.44%)
=3209.54 million